OpenStudy (anonymous):

A pizza with a diameter of 12 inches is cut into 8 equal slices. To the nearest square inch, what is the area of each slice of the pizza?

OpenStudy (anonymous):

Area = Pi x 6^2 = 36 Pi Area of 1 slice = 36 Pi / 8 = 9/2 Pi = 14.14 (2dp)

OpenStudy (anonymous):

@Zoe9639 no direct answer against the code of conduct